OpenStudy (anonymous):

Which is the equation of a line parallel to the line with equation y = 3x - 4?

OpenStudy (anonymous):

The one with slope =3.

OpenStudy (anonymous):

x - 3y = 1 3x + y = 7 x + 3y = 4 3x - y = -2

OpenStudy (anonymous):

3x - y = -2 y= 3 x +2